# Instantaneous Polarimetry with Zak-OTFS

### v1

@@ Abstract
Polarimetry, which is the ability to measure the scattering response of the environment across orthogonal polarizations, is fundamental to enhancing wireless communication and radar system performance. In this paper, we utilize the Zak-OTFS modulation to enable instantaneous polarimetry within a single transmission frame. We transmit a Zak-OTFS carrier waveform and a spread carrier waveform mutually unbiased to it simultaneously over orthogonal polarizations. The mutual unbiasedness of the two waveforms enables the receiver to estimate the full polarimetric response of the scattering environment from a single received frame. Unlike existing methods for instantaneous polarimetry with computational complexity quadratic in the time-bandwidth product, the proposed method enables instantaneous polarimetry at complexity that is only sublinear in the time-bandwidth product. Via numerical simulations, we show ideal polarimetric target detection and parameter estimation results with the proposed method, with improvements in performance and computational complexity over comparable baselines.
Index Terms:
6G, Polarimetry, Integrated Sensing and Communication, Zak-OTFS
@@ I Introduction
Polarimetry is enabled in radar and communication systems by transmitting and receiving on two orthogonal polarizations, e.g., on vertical and horizontal polarizations. The receiver estimates the $2\times 2$ polarimetric scattering response of the wireless/radar channel across all four combinations of transmit and receive polarizations. A standard approach is to transmit polarized waveforms sequentially across two frames [11, 12, 13, 14, 15]; see Fig. 1(1(a)) for an example with frequency modulated continuous wave (FMCW) transmissions. From its measurements in each frame, the receiver estimates $2\times 1$ slices of the full $2\times 2$ polarimetric scattering response. Such an approach does not provide instantaneous estimates of the scattering response within a single frame. Changes in the scattering environment between the two frames (due to mobility) may partially decorrelate the obtained estimates [12, 13, 14, 15, 16]. Sequential polarimetry also prevents frame-by-frame processing & increases the system latency, which is a critical factor for radar and communication performance in highly dynamic environments. When utilizing continuous waveforms such as FMCW (which is the waveform of choice across radar systems [19, 20, 21, 22]), the computational complexity of sequential polarimetry is quadratic in the time-bandwidth product [23, 24, 25].
@@
To unlock the full benefits of polarimetry, it is crucial to estimate the $2\times 2$ polarimetric scattering response instantaneously within a single transmission frame. Previous work [13, 14, 15] has proposed transmitting mutually unbiased waveforms, i.e., waveforms with small inner products, simultaneously across orthogonal polarizations. Mutual unbiasedness ensures that the contribution of the other waveform looks like noise to the receiver when it projects its measurements onto the basis of one of the transmit waveforms. Projecting its measurements onto the basis of each of the two transmit waveforms provides the receiver with an estimate of different $2\times 1$ slices of the polarimetric scattering response; thus enabling full $2\times 2$ polarimetric scattering response estimation from a single received frame. Mutually unbiased waveforms have been designed in prior work [13, 14, 15] via phase-coding, i.e., by modulating a common carrier waveform, e.g., a rectangular waveform, with mutually unbiased sequences, e.g., Zadoff-Chu sequences with distinct roots [13] or complementary Golay pairs [14, 15]. While this approach offers excellent polarimetric target detection and parameter estimation (detailed later in Section IV), the computational complexity of polarimetry via phase-coding remains quadratic in the time-bandwidth product [25, 26]. Moreover, we have shown in [25] that separate selection of sequences and carrier waveforms, as done in phase-coding, may be sub-optimal from a radar waveform design perspective.
@@
In this paper, we take an alternate approach to constructing mutually unbiased waveforms for instantaneous polarimetry by utilizing the Zak-OTFS (orthogonal time frequency space) modulation [27, 28, 29] in place of phase-coding. The carrier waveform in Zak-OTFS, termed pulsone, is a pulse localized in the delay-Doppler domain. We utilize a generalized discrete affine Fourier transform [25, 30] to transform the pulsone into a spread waveform that is mutually unbiased to the pulsone. Our approach has two advantages over phase-coding. First, scattering response estimation with pulsones and their unitary transformations is possible with computational complexity only sublinear in the time-bandwidth product [25]. Second, unlike phase-coding, Zak-OTFS enables joint optimization of sequences and carrier waveforms, which has been shown to be optimal from a radar waveform design perspective in [25]. We illustrate our proposed approach in Fig. 1(1(b)) and describe it in more detail in Section III.
Table I places our contributions in the context of prior work, which is described in more detail in Section II. In addition to the advantages in computational complexity and latency, our proposed approach also improves upon the number of simultaneously detectable targets and the Doppler resolution over the sequential approach based on FMCW. Moreover, in Section IV we demonstrate that the proposed approach achieves ideal polarimetric target detection and parameter estimation [^1] The results in this paper are limited to polarimetric target detection and parameter estimation in the presence of noise. Extensions to clutter [21, 31, 32, 33, 34, 10] are possible, but are not pursued in this paper. similar to the phase-coded approach at a considerably smaller computational complexity.
@@ III-A Overview of Zak-OTFS
The Zak-OTFS carrier waveform is a pulse in the delay-Doppler (DD) domain, formally a quasi-periodic localized function termed the DD pulsone. The DD pulsone is characterized by a delay period $\tau_{p}$ and a Doppler period $\nu_{p}$, with $\tau_{p}\nu_{p}=1$. The DD pulsone occupies infinite time and bandwidth. For practical implementation, the DD pulsone is limited to a time interval $T$ and a bandwidth $B$ via DD domain pulse shaping. The DD pulsone defines an orthonormal basis within the delay and Doppler periods with $BT=MN$ basis elements at $M=\nicefrac{{\tau_{p}}}{{\nicefrac{{1}}{{B}}}}=B\tau_{p}$ distinct locations along delay and $N=\nicefrac{{\nu_{p}}}{{\nicefrac{{1}}{{T}}}}=T\nu_{p}$ distinct locations along Doppler.
The discrete time domain (TD) representation of a DD pulsone on converting the DD signal to time via the inverse Zak transform followed by Nyquist sampling is [27, 28, 29]:
@@
where $\mathbf{h}_{\mathsf{eff}}[k,l]$ denotes the effective channel [^4] The effective channel approximates the physical channel when all paths are resolvable in delay with bandwidth $B$ and in Doppler with time $T$. that encompasses the effects of the physical scattering environment and transmit & receive pulse shaping/matched filtering [29, Eq. (7)], and $\mathbf{w}[n]$ denotes the additive noise at the receiver. For a scattering environment with $P$ paths/targets, let $\mathbf{h}_{\mathrm{phy}}(\tau,\nu)=\sum_{t=1}^{P}h^{(p)}\delta(\tau-\tau_{t})\delta(\nu-\nu_{t})$ denote the corresponding channel representation in the continuous DD domain. The effective channel is given by samples of the continuous effective channel, $\mathbf{h}_{\mathrm{eff}}[k,l]=\mathbf{h}_{\mathrm{eff}}\big{(}\tau=\frac{k\tau_{p}}{M},\nu=\frac{l\nu_{p}}{N}\big{)}$, where [27, 28, 29]:
@@
Accurate channel estimation is possible when the sequence $\mathbf{x}$ satisfies the crystallization condition [27, 28, 29, 35, 30]. Let $\mathcal{S}=\big{\{}(k,l)\big{|}\big{|}\mathbf{A}_{\mathbf{x},\mathbf{x}}[k,l]\big{|}=1\big{\}}$ denote the DD locations where the self-ambiguity function of $\mathbf{x}$ is unimodular, and let $\mathcal{C}$ denote the maximum DD support of the scattering environment777e.g., $\mathcal{C}=[k_{\min},k_{\max}]\times[l_{\min},l_{\max}]$ based on prior knowledge of the minimum/maximum delay and Doppler spreads of the scattering environment.. The crystallization condition requires:
@@ III-C Proposed Approach for Instantaneous Polarimetry
We enable instantaneous polarimetry at sublinear complexity by designing mutually unbiased sequences via Zak-OTFS. To that end, in the following we define the generalized discrete affine Fourier transform (GDAFT) [25, 30], which maps pulsones in (2) to mutually unbiased spread waveforms. We have previously utilized the GDAFT for designing radar waveform libraries in [25] and for spread carrier communication in [30].
@@
where $C_{(k_{0},l_{0})}[k,l]$ is a complex phase, $\big{|}C_{(k_{0},l_{0})}[k,l]\big{|}=1$, and $\mathbf{A}_{\mathbf{y},\mathbf{x}}[k,l]$ denotes the cross-ambiguity function as in (III-A). Moreover, the GDAFT preserves the sublinear computational complexity of cross-ambiguity-based channel estimation [25].
@@
For accurate channel estimation, the GDAFT parameters $A,B,C$ in Theorem 1 are chosen such that $\mathbf{c}[n]$ satisfies the crystallization condition in (7) for all four components of the polarimetric effective channel $\mathbf{h}_{\mathsf{eff}}^{(j,i)}[k,l]$, for all $i,j\in\{\mathsf{V},\mathsf{H}\}$.
@@
where $C[k,l]$ is a phase term similar to that in (9). Since the second term is the twisted convolution of the effective channel $\mathbf{h}_{\mathsf{eff}}^{(j,\bar{i})}[k,l]$ with a constant amplitude term, it simply raises the noise floor of the channel estimate. Computing each cross-ambiguity term only incurs $\mathcal{O}(BT\log T)$ complexity [25], and the overall complexity remains sublinear in $BT$.
@@ IV Numerical Results
We now qualitatively and quantitatively compare the performance of the proposed approach from Section III with sequential polarimetry via FMCW (Section II-A) and instantaneous polarimetry via phase-coding (Section II-B). We also compare against uni-polarized systems. As mentioned in the Introduction, our results are limited to polarimetry in noise; extensions to clutter are a subject of future work.
@@ IV-A Polarimetric Channel Estimation (Qualitative)
Fig. 2 qualitatively compares the performance of polarimetric channel estimation via the cross-ambiguity function in (III-A) for a four-target scattering environment. We simulate two targets with equal $h^{(p)}_{\mathsf{HH}}=0.7$ & $h^{(p)}_{\mathsf{HV}}=h^{(p)}_{\mathsf{VH}}=h^{(p)}_{\mathsf{VV}}=0$, and two targets with unequal $h^{(p)}_{\mathsf{HV}}=h^{(p)}_{\mathsf{VH}}\in\{0.3,0.95\}$ & $h^{(p)}_{\mathsf{HH}}=h^{(p)}_{\mathsf{VV}}=0$. We observe that polarimetry via FMCW detects false targets in the $\mathsf{HH}$ channel and fails to detect the low energy target in the $\mathsf{VH}$ channel due to high sidelobes of the waveform. In contrast, polarimetry via phase-coding and Zak-OTFS achieves ideal target detection with minimal sidelobes around the target locations.
@@ IV-B2 Parameter Estimation
To fuse the target parameter estimates obtained across multiple polarimetric components, we compute their weighted average using entropy-based weights, $w^{(j,i)}=1-\nicefrac{{\mathcal{H}\big{(}\widehat{\mathbf{h}}_{\mathsf{eff}}^{(j,i)}\big{)}}}{{\log_{2}(MN)}}$ for all $i,j\in\{\mathsf{V},\mathsf{H}\}$, where $\mathcal{H}(\mathbf{h})=-\sum_{k,l}\big{(}\nicefrac{{|\mathbf{h}[k,l]|^{2}}}{{\sum_{k^{\prime},l^{\prime}}|\mathbf{h}[k^{\prime},l^{\prime}]|^{2}}}\big{)}\log_{2}\big{(}\nicefrac{{|\mathbf{h}[k,l]|^{2}}}{{\sum_{k^{\prime},l^{\prime}}|\mathbf{h}[k^{\prime},l^{\prime}]|^{2}}}\big{)}$ denotes the entropy [36] of a DD channel $\mathbf{h}$. Intuitively, such entropy-based weighting prioritizes parameter estimates obtained from polarimetric components with large variation in channel amplitudes (indicating the presence of targets) as opposed to those with little variation in channel amplitudes (indicating noise and the absence of any target). For illustration, consider the extreme case with no targets. Due to mutual unbiasedness, the estimated channel per (III-C) in this case has a constant energy level of $\nicefrac{{1}}{{MN}}$ and an entropy of $\log_{2}(MN)$, for which the chosen weight is $w^{(j,i)}=0$, i.e., we do not allow parameter estimates obtained from such polarimetric components to bias our estimates.
Figs. 4(4(b))-(4(c)) plot the root-mean-squared error (RMSE) for delay and Doppler estimation for all considered systems. The RMSE for delay estimation is similar for polarimetry with FMCW, phase-coding and Zak-OTFS at high SNRs. The Doppler RMSE matches for polarimetry with phase-coding and Zak-OTFS, with $\sim 1.5\times$ improvement over FMCW, consistent with the explanation in Section II-A. Uni-polarized systems have significantly poorer delay and Doppler RMSE. Note that the RMSEs do not improve beyond a certain threshold due to the inherent resolution limits of the chosen waveforms and the parameter estimation procedure described previously. Designing optimal parameter estimation approaches that achieve information theoretic lower bounds, e.g., the Cramér-Rao bound [37], is a subject of future work.

### v3

@@ Abstract
Polarimetry, which is the ability to measure the scattering response of the environment across orthogonal polarizations, is fundamental to enhancing wireless communication and radar system performance. In this paper, we utilize the Zak-OTFS modulation to enable instantaneous polarimetry within a single transmission frame. We transmit a Zak-OTFS carrier waveform and a spread carrier waveform mutually unbiased to it simultaneously over orthogonal polarizations. The mutual unbiasedness of the two waveforms enables the receiver to estimate the full polarimetric response of the scattering environment from a single received frame. Unlike existing methods for instantaneous polarimetry with computational complexity quadratic in the time-bandwidth product, the proposed method enables instantaneous polarimetry at near-linear complexity in the time-bandwidth product. Via numerical simulations, we show ideal polarimetric target detection and parameter estimation results with the proposed method, with improvements in computational complexity and greater clutter resilience over comparable baselines.
@@ I Introduction
Polarimetry is enabled in radar and communication systems by transmitting and receiving on two orthogonal polarizations, e.g., on vertical and horizontal polarizations. The receiver estimates the $2\times 2$ polarimetric scattering response of the wireless/radar channel across all four combinations of transmit and receive polarizations. A standard approach is to transmit polarized waveforms sequentially across two frames [11, 12, 13, 14, 15]; see Fig. 1(1(a)) for an example with frequency modulated continuous wave (FMCW) transmissions. From its measurements in each frame, the receiver estimates $2\times 1$ slices of the full $2\times 2$ polarimetric scattering response. Such an approach does not provide instantaneous estimates of the scattering response within a single frame. Changes in the scattering environment between the two frames (due to mobility) may partially decorrelate the obtained estimates [12, 13, 14, 15, 16]. Sequential polarimetry also prevents frame-by-frame processing & increases the system latency, which is a critical factor for radar and communication performance in highly dynamic environments. When utilizing continuous waveforms, such as FMCW and pulsed waveforms [19, 20, 21, 22], the computational complexity of sequential polarimetry is quadratic in the time-bandwidth product [23, 24, 25].
(a) Sequential Polarimetry with FMCW
(b) Instantaneous Polarimetry with Zak-OTFS (Ours)
@@
To unlock the full benefits of polarimetry, it is crucial to estimate the $2\times 2$ polarimetric scattering response instantaneously within a single transmission frame. Previous work [13, 14, 15] has proposed transmitting mutually unbiased waveforms [^1] The term “mutually unbiased” is from quantum information theory [26]. Formally, two $d$-length waveforms are mutually unbiased if their inner product has magnitude $\nicefrac{{1}}{{\sqrt{d}}}$. Measurements from one waveform are “statistically independent” to those from the other waveform with uniform probability $\nicefrac{{1}}{{d}}$., i.e., waveforms with small inner products, simultaneously across orthogonal polarizations. Mutual unbiasedness ensures that the contribution of the other waveform looks like noise to the receiver when it projects its measurements onto the basis of one of the transmit waveforms. Projecting its measurements onto the basis of each of the two transmit waveforms provides the receiver with an estimate of different $2\times 1$ slices of the polarimetric scattering response; thus enabling full $2\times 2$ polarimetric scattering response estimation from a single received frame. Mutually unbiased waveforms have been designed in prior work [13, 14, 15] via phase-coding, i.e., by modulating a common carrier waveform, e.g., a rectangular waveform, with mutually unbiased sequences, e.g., Zadoff-Chu sequences with distinct roots [13] or complementary Golay pairs [14, 15]. While this approach offers excellent polarimetric target detection and parameter estimation (detailed later in Section IV), the computational complexity of polarimetry via phase-coding remains quadratic in the time-bandwidth product [25, 27]. Moreover, we have shown in [25] that separate selection of sequences and carrier waveforms, as done in phase-coding, may be sub-optimal from a radar waveform design perspective.
@@
In this paper, we take an alternate approach to constructing mutually unbiased waveforms for instantaneous polarimetry by utilizing the Zak-OTFS (orthogonal time frequency space) modulation [28, 29, 30] in place of phase-coding. The carrier waveform in Zak-OTFS, termed pulsone, is a pulse localized in the delay-Doppler domain. We utilize a generalized discrete affine Fourier transform [25, 31] to transform the pulsone into a spread waveform that is mutually unbiased to the pulsone. Our approach has two advantages over phase-coding. First, scattering response estimation with pulsones and their unitary transformations is possible with near-linear computational complexity in the time-bandwidth product [25]. Second, unlike phase-coding, Zak-OTFS enables joint optimization of sequences and carrier waveforms, which has been shown to be optimal from a radar waveform design perspective in [25]. We illustrate our proposed approach in Fig. 1(1(b)) and describe it in more detail in Section III.
Table I places our contributions in the context of prior work, which is described in more detail in Section II. In addition to the advantages in computational complexity and latency, our proposed approach also improves upon the number of simultaneously detectable targets and the Doppler resolution over the sequential approach based on FMCW. Moreover, in Section IV we demonstrate that the proposed approach achieves ideal polarimetric target detection and parameter estimation [^2] The results in this paper serve as a proof-of-concept numerical demonstration of the concept; hardware demonstration will be pursued in future work. Potential challenges include managing the peak-to-average power ratio (PAPR) and over-the-air synchronization. With regards to the former, we note that a natural advantage of the proposed framework is its low PAPR of only $5.6$ dB; see [25, Section 4.2] and [31] for details. This reduction in PAPR has also been experimentally demonstrated in [32] in the sub-terahertz band. and greater clutter resilience compared to the phase-coded approach at smaller computational complexity.
@@ III-A Overview of Zak-OTFS
The Zak-OTFS carrier waveform is a pulse in the delay-Doppler (DD) domain, formally a quasi-periodic localized function termed the DD pulsone [^5] Termed “pulsone” due to its structure of a pulse train modulated by a tone in the time domain, see [29, Fig. 2] for an illustration.. The DD pulsone is characterized by a delay period $\tau_{p}$ and a Doppler period $\nu_{p}$, with $\tau_{p}\nu_{p}=1$. The DD pulsone occupies infinite time and bandwidth. For practical implementation, the DD pulsone is limited to a time interval $T$ and a bandwidth $B$ via DD domain pulse shaping. The DD pulsone defines an orthonormal basis within the delay and Doppler periods with $BT=MN$ basis elements at $M=\nicefrac{{\tau_{p}}}{{\nicefrac{{1}}{{B}}}}=B\tau_{p}$ distinct locations along delay and $N=\nicefrac{{\nu_{p}}}{{\nicefrac{{1}}{{T}}}}=T\nu_{p}$ distinct locations along Doppler.
The DD pulsone is converted to a time domain (TD) waveform via the inverse Zak transform [28, 29, 30]. After Nyquist sampling, the discrete TD pulsone waveform is [28, 29, 30]:
@@
where $\mathbf{h}_{\mathsf{eff}}[k,l]$ denotes the effective channel [^6] The effective channel approximates the physical channel when all paths are resolvable in delay with bandwidth $B$ and in Doppler with time $T$. that encompasses the effects of the physical scattering environment and transmit & receive pulse shaping/matched filtering [30, Eq. (7)], and $\mathbf{w}[n]$ denotes the additive noise at the receiver. For a scattering environment with $P$ paths/targets, let $\mathbf{h}_{\mathrm{phy}}(\tau,\nu)=\sum_{t=1}^{P}h^{(p)}\delta(\tau-\tau_{t})\delta(\nu-\nu_{t})$ denote the corresponding channel representation in the continuous DD domain. The effective channel is given by samples of the continuous effective channel, $\mathbf{h}_{\mathrm{eff}}[k,l]=\mathbf{h}_{\mathrm{eff}}\big(\tau=\frac{k\tau_{p}}{M},\nu=\frac{l\nu_{p}}{N}\big)$, where [28, 29, 30]:
@@
Accurate channel estimation is possible when the sequence $\mathbf{x}$ satisfies the crystallization condition [28, 29, 30, 33, 31]. Let $\mathcal{S}=\big\{(k,l)\big|\big|\mathbf{A}_{\mathbf{x},\mathbf{x}}[k,l]\big|=1\big\}$ denote the DD locations where the self-ambiguity function of $\mathbf{x}$ is unimodular, and let $\mathcal{C}$ denote the maximum DD support of the scattering environment999e.g., $\mathcal{C}=[k_{\min},k_{\max}]\times[l_{\min},l_{\max}]$ based on prior knowledge of the minimum/maximum delay and Doppler spreads of the scattering environment.. The crystallization condition requires:
@@ III-C Proposed Approach for Instantaneous Polarimetry
We enable instantaneous polarimetry at near-linear complexity by designing mutually unbiased sequences via Zak-OTFS. To that end, in the following we define the generalized discrete affine Fourier transform (GDAFT) [25, 31], which maps pulsones in (2) to mutually unbiased spread waveforms. We have utilized the GDAFT to design radar waveform libraries in [25] and for spread carrier communication in [31].
@@
where $C_{(k_{0},l_{0})}[k,l]$ is a complex phase, $\big|C_{(k_{0},l_{0})}[k,l]\big|=1$, and $\mathbf{A}_{\mathbf{y},\mathbf{x}}[k,l]$ denotes the cross-ambiguity function as in (III-A). Moreover, the GDAFT preserves the near-linear computational complexity of cross-ambiguity-based channel estimation [25].
@@
For accurate channel estimation, the GDAFT parameters $A,B,C$ in Theorem 1 are chosen such that $\mathbf{c}[n]$ satisfies the crystallization condition in (7) for all four components of the polarimetric effective channel $\mathbf{h}_{\mathsf{eff}}^{(j,i)}[k,l]$, for all $i,j\in\{\mathsf{V},\mathsf{H}\}$. For a detailed discussion on GDAFT parameter selection, see Fig. 1 and the associated example in [31, Section IV-C].
@@
where $C[k,l]$ is a phase term similar to that in (9). Since the second term is the twisted convolution of the effective channel $\mathbf{h}_{\mathsf{eff}}^{(j,\bar{i})}[k,l]$ with a constant amplitude term, it simply raises the noise floor of the channel estimate. Computing each cross-ambiguity term only incurs $\mathcal{O}(BT\log T)$ complexity [^10] The complexity reduction from quadratic to near-linear is due to the symmetry of the Zak-OTFS carrier waveform. Specifically, the cross-ambiguity function computation with Zak-OTFS carrier waveforms reduces to an FFT (fast Fourier transform) calculation. For more details, see [25, Section 4.3]., and the overall complexity remains near-linear in $BT$.
@@ IV Numerical Results
We now qualitatively and quantitatively compare the performance of the proposed approach from Section III with sequential polarimetry via FMCW (Section II-A) and instantaneous polarimetry via phase-coding (Section II-B). We also compare against uni-polarized systems. Our initial results are limited to target detection and estimation in the presence of noise; extensions to clutter are briefly pursued in Section IV-C.
@@ IV-A Polarimetric Channel Estimation (Qualitative)
[…]
@@ IV-B2 Parameter Estimation
To fuse the target parameter estimates obtained across multiple polarimetric components, we compute their weighted average using entropy-based weights, $w^{(j,i)}=1-\nicefrac{{\mathcal{H}\big(\widehat{\mathbf{h}}_{\mathsf{eff}}^{(j,i)}\big)}}{{\log_{2}(MN)}}$ for all $i,j\in\{\mathsf{V},\mathsf{H}\}$, where $\mathcal{H}(\mathbf{h})=-\sum_{k,l}\big(\nicefrac{{|\mathbf{h}[k,l]|^{2}}}{{\sum_{k^{\prime},l^{\prime}}|\mathbf{h}[k^{\prime},l^{\prime}]|^{2}}}\big)\log_{2}\big(\nicefrac{{|\mathbf{h}[k,l]|^{2}}}{{\sum_{k^{\prime},l^{\prime}}|\mathbf{h}[k^{\prime},l^{\prime}]|^{2}}}\big)$ denotes the entropy [34] of a DD channel $\mathbf{h}$. Intuitively, such entropy-based weighting prioritizes parameter estimates obtained from polarimetric components with large variation in channel amplitudes (indicating the presence of targets) as opposed to those with little variation in channel amplitudes (indicating noise and the absence of any target). For illustration, consider the extreme case with no targets. Due to mutual unbiasedness, the estimated channel per (III-C) in this case has a constant energy level of $\nicefrac{{1}}{{MN}}$ and an entropy of $\log_{2}(MN)$, for which the chosen weight is $w^{(j,i)}=0$, i.e., such components are not allowed to bias our estimates.
Figs. 4(4(b))-(4(c)) plot the root-mean-squared error (RMSE) for delay and Doppler estimation for all considered systems. The RMSE for delay estimation is similar for polarimetry with FMCW, phase-coding and Zak-OTFS at high SNRs. The Doppler RMSE matches for polarimetry with phase-coding and Zak-OTFS, with $\sim 1.5\times$ improvement over FMCW, consistent with the explanation in Section II-A. Uni-polarized systems have significantly poorer delay and Doppler RMSE. Note that the RMSEs do not improve beyond a certain threshold due to the inherent resolution limits of the chosen waveforms. Future work will design optimal approaches achieving theoretical bounds, e.g., the Cramér-Rao bound [35].
IV-C Target Detection in Constant-$\gamma$ Clutter
Fig. 5 plots the detection ROC curve for single-target detection in constant-$\gamma$ clutter for all dual-polarized systems. For a metropolitan terrain and a carrier frequency of $f_{c}=4$ GHz, we obtain the clutter parameter value $\gamma=-1.99$ dB from [36]. Fig. 5 shows that instantaneous polarimetry via Zak-OTFS exhibits greater resilience to clutter as compared to methods based on phase-coded and FMCW transmissions.
Figure 5: Instantaneous polarimetry via Zak-OTFS exhibits greater resilience to clutter as compared to competing methods.